# Guided filtering and smoothing for infinite-dimensional diffusions

### v1

@@ Abstract.
We consider the filtering and smoothing problems for an infinite-dimensional diffusion process $X$, observed through a finite-dimensional representation at discrete points in time. At the heart of our proposed methodology lies the construction of a path measure, termed the guided distribution of $X$, that is absolutely continuous with respect to the law of $X$, conditioned on the observations. We show that this distribution can be incorporated as a potent proposal measure for both sequential Monte Carlo as well as Markov Chain Monte Carlo schemes to tackle the filtering and smoothing problems respectively. In the offline setting, we extend our approach to incorporate parameter estimation of unknown model parameters. The proposed methodology is numerically illustrated in a case study for the stochastic Amari equation.
@@ 1. Introduction
Consider an infinite-dimensional diffusion equation of the form
@@
Here, $A$ denotes the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ on a Hilbert space $H$, whereas $F$ is a Lipschitz continuous non-linearity and $Q$ a symmetric, positive operator of trace class on $H$. The process $(W_{t})_{t\geq 0}$ is a cylindrical Wiener process on $H$, defined on a stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ and $\mu_{0}$ denotes a Borel measure on $H$. Infinite-dimensional diffusions like that of Equation 1 arise, for example, in the context of semilinear stochastic partial differential equations or stochastic field equations.
Throughout this work, we assume that Equation 1 admits a unique mild solution $X$ satisfying
@@
for any Borel set $B\subset H$. Moreover, if Equation 1 is parametrised by some unknown model parameters $\theta\in\mathbb{R}^{p}$, we are interested in parameter estimation of $\theta$ in the offline setting.
@@ 1.1. Approach
Our approach relies heavily on the exponential changes of measure studied in [Pieper-Sethmacher et al., 2025a]. Consider for now the case that $n=1$, i.e. we observe one realisation $y$ of the random variable $Y\mid X_{T}\sim k(X_{T},\cdot)$. In this simplified case, estimating the conditional distribution of $X$ given $Y=y$ solves both the filtering as well as the smoothing problem.
@@
for any bounded and measurable function $\varphi$. Moreover, under certain regularity conditions on $h$, it was shown in [Pieper-Sethmacher et al., 2025a] that the change of measure $\mathbb{P}^{h}$ is of Girsanov type. The process $X$ under $\mathbb{P}^{h}$ is then a mild solution to the SPDE
@@
    (8) $\mathop{}\!\mathrm{d}X^{g}_{t}$ $=\left[AX^{g}_{t}+F(t,X^{g}_{t})+Q\mathop{}\!\mathrm{D}_{x}\log g(t,X^{g}_{t})\right]\mathop{}\!\mathrm{d}t+Q^{\frac{1}{2}}\mathop{}\!\mathrm{d}W_{t},$
@@
The function $x\mapsto g(t,x)$ is to be interpreted as the likelihood function of $Z_{t}=x$ under the simplified dynamics of $Z$ and therefore satisfies condition (i). Moreover, condition (ii) is ensured by the Gaussian nature of $g(t,x)$ implying a global Lipschitz condition of the score function $\mathop{}\!\mathrm{D}_{x}\log g(t,x)$. Lastly, by construction of $\mathbb{P}^{h}$ and $\mathbb{P}^{g}$, the laws of $X^{g}$ and $X^{h}$ are absolutely continuous in path space, hence satisfying the third condition imposed on $g$. In [Pieper-Sethmacher et al., 2025b] it was shown that this absolute continuity persists if the observation kernel degenerates to a Dirac measure in $Lx$, i.e. if $Y\mid X_{T}\sim\delta_{Lx}(\cdot)$. In other words, the guided distribution remains a valid importance sampling distribution for $X^{h}$ even under highly informative observations.
Our aim in this paper is to apply this basic idea of constructing the conditioned process $X^{h}$ as well as the guided process $X^{g}$ by a change of measure to derive solutions to both the filtering as well as the smoothing problem.
For the filtering problem, we propose the law of $X^{g}$ as a natural candidate for the proposal distribution in a particle filter (or sequential Monte Carlo) approach. The particle filter is a well-established algorithm that approximates the filtering distribution in (5) by a set of samples (‘particles’) $X_{t_{i}}^{(j)}$ with associated weights $w_{i}^{(j)}$. At each iteration $i$, the particles are evolved forward following a proposal distribution $X_{[t_{i},t_{i+1}]}^{(j)}\sim\mathbb{Q}(\cdot\mid X_{t_{i}}^{(j)})$ with the associated weights following the recursive update
$w_{i+1}^{(j)}\propto w_{i}^{(j)}\,k(X^{(j)}_{t_{i+1}},y_{i+1})\frac{\mathop{}\!\mathrm{d}\mathbb{P}\left(X^{(j)}_{[t_{i},t_{i+1}]}\in\cdot\mid X_{t_{i}}^{(j)}\right)}{\mathop{}\!\mathrm{d}\mathbb{Q}(\cdot\mid X_{t_{i}}^{(j)})}\left(X^{(j)}_{[t_{i},t_{i+1}]}\right).$
It is well-known that the locally optimal proposal distribution $\mathbb{Q}$ is given by the conditional distribution of $X^{(j)}_{[t_{i},t_{i+1}]}\mid X^{(j)}_{t_{i}},Y_{i+1}=y_{i+1}$ (cf. 10.3, [Chopin et al., 2020]). This corresponds to the $h$-transform introduced above with the intractable function
$h_{i}(t,x)$ $=\int_{H}k(z,y_{i+1})\mu_{t,t_{i+1}}(x,\mathop{}\!\mathrm{d}z),\quad t\in(t_{i},t_{i+1}],x\in H.$
A natural candidate for a substitute proposal distribution is therefore the guided distribution corresponding to the observation $y_{i+1}$.
On the other hand, the smoothing problem can be solved after extending the change of measure $\mathbb{P}^{h}$ to include the complete set of observations $(Y_{1},\ldots,Y_{n})=(y_{1},\ldots,y_{n})$ by defining the function
$h(t,x):=\int k(x_{t_{i}},y_{i})\left(\prod_{j=i}^{n-1}k(x_{t_{j+1}},y_{j+1})\,\mu_{t_{j},t_{j+1}}(x_{t_{j}},\mathop{}\!\mathrm{d}x_{t_{j+1}})\right)\mu_{t,t_{i}}(x,\mathop{}\!\mathrm{d}x_{t_{i}}),\quad t\in(t_{i-1},t_{i}].$
In Section 2 we show that $h$ defines a change measure $\mathbb{P}^{h}$ under which the law of $X^{h}$ is exactly the smoothing distribution (6) of $X$. Substituting again the Markov kernel $\mu$ with the kernel $\nu$ of an auxiliary process $Z$ leads to a tractable change of measure $\mathbb{P}^{g}$ and a corresponding guided process $X^{g}$ that satisfies yet another SPDE, while maintaining absolute continuity of the laws of $X^{h}$ and $X^{g}$ with a tractable Radon-Nikodym derivative $\Phi$.
For certain choices of $Z$, we show that $g$ can be directly computed on the basis of the recursion known as the backwards information filter
(9) $\begin{split}g(t,x)&=\int g(t_{i},x_{t_{i}})\,\nu_{t,t_{i}}(x,\mathop{}\!\mathrm{d}x_{t_{i}}),\quad t\in(t_{i-1},t_{i}],\\ g(t_{i},x)&=k(x,y_{j})\,g^{+}(t_{i},x),\quad i\in\{1,\dots,n\},\end{split}$
with $g^{+}(t_{i},x):=\lim_{\begin{subarray}{c}t\downarrow t_{i}\\ t\neq t_{i}\end{subarray}}g(t,x)$. This enables drawing weighted samples of $X^{h}$ which can then be passed onto general Markov chain Monte Carlo or importance sampling schemes to sample from the smoothing distribution of $X$.
@@ 2. Changes of measure
    the distribution $\mathcal{L}^{g}(X)$ can be sampled by numerically solving the SPDE
@@ 2.1. The smoothing distribution
    In addition, if $h$ is Fréchet differentiable in $x$ such that $\mathop{}\!\mathrm{D}_{x}h\in C_{m}((t_{i-1},t_{i});H)$ for all $i=1\ldots n$, then the process $X$ under $\mathbb{P}^{h}$ is a mild solution to the SPDE
@@ 2.2. The guided distribution
For this, let $Z$ be the Ornstein-Uhlenbeck process, defined as the unique mild solution to the SPDE
@@
and let $R_{t}\in L(\mathbb{R}^{m^{+}_{i}})$ and $\alpha_{t}\in\mathbb{R}^{m^{+}_{i}}$ be defined by the backwards recursions
@@
    Under the measure $\mathbb{P}^{g}$ defined on $\mathcal{F}_{T}$ by $\mathop{}\!\mathrm{d}\mathbb{P}^{g}_{\mid\mathcal{F}_{T}}=E^{g}_{T}\mathop{}\!\mathrm{d}\mathbb{P}_{\mid\mathcal{F}_{T}}$, the process $X$ is the unique mild solution to the SPDE
@@
Hence, in total, the cost of the operations required to compute $G$ using Theorem 2.6 on a spatio-temporal grid with $(N,M)$ grid points is of complexity
@@ 2.3. Computationally efficient guiding
(29) $\begin{split}\begin{cases}\mathop{}\!\mathrm{d}U_{t}&=\left[-A^{*}U_{t}-U_{t}A+U_{t}QU_{t}\right]\mathop{}\!\mathrm{d}t\\ U_{t_{i}}&=L^{*}\Sigma^{-1}L+U^{+}_{t_{i}},\end{cases}\end{split}$
@@
(30) $\begin{split}\begin{cases}\mathop{}\!\mathrm{d}V_{t}&=\left[-A^{*}V_{t}+U_{t}QV_{t}+U_{t}a_{t}\right]\mathop{}\!\mathrm{d}t,\\ V_{t_{i}}&=L^{*}\Sigma^{-1}y_{i}+V^{+}_{t_{i}}.\end{cases}\end{split}$
@@
Without going into too much detail, it is worth pointing out that there exists extensive literature on the approximation of the infinite-dimensional Ricatti equation. We refer to [Gibson, 1983], [Burns and Rautenberg, 2015], [Cheung, 2025] and references within for an introduction into the relevant literature.
@@
which are solved in closed form for all $j=1\ldots m$ by (cf. [Zaitsev and Polyanin, 2002], Chapter 1.1)
@@
(31) $\begin{split}\begin{cases}\mathop{}\!\mathrm{d}c_{t}&=\left[\dfrac{1}{2}\operatorname{tr}\left[U_{t}Q\right]-\left\langle a_{t},V_{t}\right\rangle-\dfrac{1}{2}\left\langle V_{t},QV_{t}\right\rangle\right]\mathop{}\!\mathrm{d}t,\\ c_{t_{i}}&=\log f(y_{i};0,\Sigma)-c^{+}_{t_{i}}.\end{cases}\end{split}$
@@ 3. Guided particle filtering
In this section, we show how a special case of the guided distribution can be used as proposal distribution in a particle filtering scheme. For the convenience of the reader, we give a brief outline of the particle filter (PF). We refer to [Doucet et al., 2001] or [Chopin et al., 2020] for book-length treatments of this field.
@@ 3. Resample: if ESS $<J_{0}$ then
The proposal $\mathbb{Q}$ minimises the variance of the weights $w_{i}^{(j)}$, see for example [Chopin et al., 2020], Theorem 10.1 for details. However, as noted in the introduction and previous section, this proposal kernel is typically intractable.
We propose instead to steer the particles on the interval $[t_{i-1},t_{i}]$ into regions of high probability given $Y_{i}=y_{i}$ by the guided distribution introduced in Section 2.2. For this, define
@@
The measure $\mathcal{L}^{g_{i}}(X_{[t_{i-1},t_{i}]})$ is the ‘one-step-ahead’ guided distribution on $[t_{i-1},t_{i}]$ corresponding to the next observation $y_{i}$. Hence, the functions $g_{i}$ and $G_{i}$ only need to be computed as functions of $y_{i}$ and the parametrisation of Theorem 2.6 remains a computationally valid choice. Specifically, on any $[t_{i-1},t_{i}]$, $g_{i}(t,x)$ is the Gaussian likelihood $g_{i}(t,x)=f(y_{i};L_{t}x+\alpha_{t},R_{t})$ with $L_{t}=LS_{t_{i}-t}$, $\alpha_{t}=L\left(\int_{t}^{t_{i}}S_{t_{i}-s}a_{s}\,\mathop{}\!\mathrm{d}s\right)$ and $R_{t}=\Sigma+LQ_{t_{i}-t}L^{*}$, whereas $G_{i}$ is the score

### v2

@@ Abstract.
We consider the filtering and smoothing problems for an infinite-dimensional diffusion process $X$, observed through a finite-dimensional representation at discrete points in time. At the heart of our proposed methodology lies the construction of a path measure, termed the guided distribution of $X$, that is absolutely continuous with respect to the law of $X$, conditioned on the observations. We show that this distribution can be incorporated as a potent proposal measure for both sequential Monte Carlo as well as Markov Chain Monte Carlo schemes to tackle the filtering and smoothing problems respectively. In the offline setting, we extend our approach to incorporate parameter estimation of unknown model parameters. The proposed methodology is numerically illustrated in a case study for a stochastic neural field equation.
@@ 1. Introduction
Enriching model predictions of a stochastic dynamical system using partial and noisy observations is an attractive idea, with numerous applications in the physical, engineering, and biological sciences [64, 40, 49, 56, 57]. From a mathematical and statistical viewpoint, such data assimilation methods constitute inverse problems in which the unknown may be the state of the system, its model parameters, or a combination thereof. If the underlying system is spatially independent, it is governed by a stochastic differential equation (SDE) in a finite-dimensional state space, with observations pertaining the state variables at selected observation times. The mathematical foundations for these problems have been developed over the past decade, and are currently available in several texts [5, 40, 56, 57].
The present paper concerns data assimilation for systems governed by infinite-dimensional SDEs. Such systems include spatiotemporal processes modelled by stochastic partial differential equations (SPDEs) as well as spatially extended SDEs in which spatial coupling is induced from nonlocal interactions. This setting poses technical challenges in view of the dimensionality of the state space, and fewer studies in this area are currently available. We introduce a unified framework for two classical data assimilation tasks – filtering and smoothing – as well as parameter estimation for SDEs in Hilbert spaces. Our framework differs from existing methods by treating filtering and smoothing in a rigorous abstract form, taking an infinite-dimensional dynamical system approach to the data assimilation problem. This aligns with the “discretisation-free” formulation of inverse problems as advocated in [65] and implies that, contrary to approaches where the SDE is first discretised, our methods are well defined under increasing mesh refinements. We demonstrate our methodology in a case study for the stochastic Wilson–Cowan–Amari equation, estimating parameters that were previously assumed to be known in the literature.
1.1. Setup
Consider a system governed by an infinite-dimensional diffusion equation of the form
@@ 1.1. Setup
Here, $A$ denotes the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ on a Hilbert space $H$, whereas $F$ is a continuous nonlinearity and $Q$ a symmetric, positive operator of trace class on $H$. The process $(W_{t})_{t\geq 0}$ is a cylindrical Wiener process on $H$, defined on a stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ and $\mu_{0}$ denotes a Borel measure on $H$.
We assume that Equation 1 admits a unique mild solution $X$ satisfying
@@
for any Borel set $B\subset H$. Moreover, if Equation 1 is parametrised by some unknown model parameter $\theta\in\mathbb{R}^{p}$, we address parameter estimation of $\theta$ in the offline setting.
1.2. Related Work
In the context of SPDEs, there has been a growing interest in statistical inference for infinite-dimensional SDEs. As the survey [13] shows, the literature mostly concerns models admitting closed-form parameter estimators studied under increasing-information asymptotics. Additionally, results for nonlinear SPDEs are sparse. In the frequentist framework, inference of a drift parameter for the two-dimensional stochastic Navier-Stokes equation has been considered in [14] and for semi-linear SPDEs in [52], [2], and [15]. Estimation concerning the nonlinearity has been treated in the parametric case in [29] and in the non-parametric case in [28, 32].
In contrast to the frequentist statistical perspective, “results on filtering in the context of SPDEs are rather scarce” and “Bayesian inference for SPDEs is another area with few existing results” ([13], page 326). Both of these problems are commonly found in the literature for spatiotemporal statistics. Here, formulating so-called ‘physically inspired spatiotemporal models’ as solutions to linear SPDEs has been proposed as a more computationally efficient approach compared to conventional methods based on Gaussian process regression, see [43], [58], and [62]. A recent introduction to this framework can be found in Chapter 5 of [69]. However, the results in this direction are limited to linear SPDEs and do not generalise to nonlinear and non-Gaussian signals.
The few works that address data assimilation for nonlinear SPDEs mostly concern the filtering problem. In [44], a guided particle filter, including tempering and resample-move steps, has been introduced to estimate the filtering distribution of a two-dimensional stochastic Navier-Stokes signal. Moreover, a similar approach was used in [39] in the context of filtering for the stochastic rotating shallow water (SRSW) model and in [20] for the Camassa-Holm equations. Likewise, our approach to the filtering problem expands on the methodology proposed in [44].
Smoothing and Bayesian parameter inference for infinite-dimensional diffusions in the generality of our setup has, to the best of our knowledge, not been considered previously in the literature. In the special case of a single, noiseless observation and known model parameters, smoothing is reduced to sampling of an infinite-dimensional diffusion bridge. This has been investigated for strong solutions to infinite-dimensional SDEs in [71] and for mild solutions in [54].
As stated above, our paper addresses also processes that evolve according to spatially extended SDEs; these arise naturally when the dynamical system of interest is inherently nonlocal, as in models coming from neuroscience [34, 26, 9, 19], ecology [67], and solid mechanics [22, 25]. The infinite-dimensional SDE that we use as an application of our theory is a stochastic version of the Wilson–Cowan–Amari equation [70, 3], also known as neural field equation. In this model the cortex is represented as a continuum, and nonlocality manifests itself through a nonlinear, nonlocal operator collecting global neuronal activity. There are multiple reasons to target this system: (i) it is a tractable example of a parabolic, nonlocal, and non-diffusive system, which sets it apart from the other applications mentioned above, modelling nonlocal diffusion or nonlocal reaction; (ii) in spite of its simplicity, it shares the salient mathematical features of more sophisticated neuroscience models; (iii) it is capable of supporting many spatiotemporal patterns observed in healthy and pathological brains, and crucially (iv) the mathematical neuroscience community has begun studying data assimilation problems for these types of systems, prompted by available brain datasets collected in the past decades.
Wilson–Cowan–Amari models subject to stochastic forcing have been introduced more recently than their PDE counterparts, with literature focusing on their analysis [35, 36, 37, 27, 46], and numerical simulation [41, 42]. Further, inverse problems for neural fields have been studied in the context of kernel reconstructions [55, 49, 1], and more specifically in data assimilation using Kalman filters with in-vitro [61, 59] and synthetic [38] data. In contrast to the existing literature, we treat the data assimilation problem in continuous space and time, rather than relying on a spatiotemporal discretisation of the neural field model.
@@ 1.3. Approach
Our approach relies on the construction of guided processes, a tractable class of processes obtained by steering the solution $X$ based on the observed data. For SDEs with Euclidean state-spaces, the idea of guided processes goes to back to [16] and [23], the terminology being first introduced in [51]. A related technique is known as nudging in the data assimilation literature. We comment on this connection in Appendix C.
Here, we build specifically upon the work on guided processes as defined in [60], [48] and [66]. To lift this approach to the infinite-dimensional setting, we rely heavily on the exponential changes of measure studied in [53].
Consider for now the case that $n=1$, i.e. we observe one realisation $y$ of the random variable $Y\mid X_{T}\sim k(X_{T},\cdot)$. In this simplified case, estimating the conditional distribution of $X$ given $Y=y$ solves both the filtering as well as the smoothing problem.
@@
for any bounded and measurable function $\varphi$. Moreover, under certain regularity conditions on $h$, it was shown in [53] that the change of measure $\mathbb{P}^{h}$ is of Girsanov type. The process $X$ under $\mathbb{P}^{h}$ is then a mild solution to
@@
    (8) $\mathop{}\!\mathrm{d}X^{g}_{t}$ $=\left[AX^{g}_{t}+F(t,X^{g}_{t})+Q\mathop{}\!\mathrm{D}_{x}\log g(t,X^{g}_{t})\right]\mathop{}\!\mathrm{d}t+Q^{\frac{1}{2}}\mathop{}\!\mathrm{d}W^{g}_{t},$
@@
The function $x\mapsto g(t,x)$ is to be interpreted as the likelihood function of $Z_{t}=x$ under the simplified dynamics of $Z$ and therefore satisfies condition (i). Moreover, condition (ii) is ensured by the Gaussian nature of $g(t,x)$ implying a global Lipschitz condition of the score function $\mathop{}\!\mathrm{D}_{x}\log g(t,x)$. Lastly, by construction of $\mathbb{P}^{h}$ and $\mathbb{P}^{g}$, the laws of $X^{g}$ and $X^{h}$ are absolutely continuous in path space, hence satisfying the third condition imposed on $g$. In [54] it was shown that this absolute continuity persists if the observation kernel degenerates to a Dirac measure in $Lx$, i.e. if $Y\mid X_{T}\sim\delta_{Lx}(\cdot)$. In other words, the guided distribution remains a valid importance sampling distribution for $X^{h}$ even under highly informative observations.
In this work we apply this basic idea of constructing the conditioned process $X^{h}$ as well as the guided process $X^{g}$ by a change of measure to derive solutions to both the filtering as well as the smoothing problem. For the filtering problem, we propose the law of $X^{g}$ as a natural candidate for the proposal distribution in a particle filter (or sequential Monte Carlo) approach.
On the other hand, to solve the smoothing problem we extend the changes of measure $\mathbb{P}^{h}$ and $\mathbb{P}^{g}$ to account for the complete set of observations $(Y_{1},\ldots Y_{n})$. This enables drawing weighted samples of $X^{h}$ which can then be passed onto Markov chain Monte Carlo or importance sampling schemes to sample from the smoothing distribution of $X$.
@@ 2. Changes of measure
    the distribution $\mathcal{L}^{g}(X)$ can be sampled by numerically solving
@@ 2.1. The smoothing distribution
    In addition, if $h$ is Fréchet differentiable in $x$ such that $\mathop{}\!\mathrm{D}_{x}h\in C_{m}((t_{i-1},t_{i});H)$ for all $i=1\ldots n$, then the process $X$ under $\mathbb{P}^{h}$ is a mild solution to
@@ 2.2. The guided distribution
For this, let $Z$ be the Ornstein-Uhlenbeck process, defined as the unique mild solution to
@@
[…]
@@
    Under the measure $\mathbb{P}^{g}$ defined on $\mathcal{F}_{T}$ by $\mathop{}\!\mathrm{d}\mathbb{P}^{g}_{\mid\mathcal{F}_{T}}=E^{g}_{T}\mathop{}\!\mathrm{d}\mathbb{P}_{\mid\mathcal{F}_{T}}$, the process $X$ is the unique mild solution to
@@
Hence, in total, the cost of the operations required to compute $G$ using Theorem 2.6 on a spatiotemporal grid with $(N,M)$ grid points is of complexity
@@ 2.3. Computationally efficient guiding
[…]
@@
[…]
@@
Without going into too much detail, it is worth pointing out that there exists extensive literature on the approximation of the infinite-dimensional Ricatti equation. We refer to [30], [10], [11] and references within for an introduction into the relevant literature.
@@
which are solved in closed form for all $j=1\ldots m$ by (cf. [72], Chapter 1.1)
@@
[…]
@@ 3. Guided particle filtering
In this section, we show how a special case of the guided distribution can be used as proposal distribution in a particle filtering scheme. For the convenience of the reader, we give a brief outline of the particle filter (PF). We refer to [24] or [12] for book-length treatments of this field.
@@ 3. Resample: if ESS $J_{0}$ then
The proposal $\mathbb{Q}$ minimises the variance of the weights $w_{i}^{(j)}$, see for example [12], Theorem 10.1 for details. However, as noted in the introduction and previous section, this proposal kernel is typically intractable.
We propose instead to steer the particles on the interval $t_{i-1},t_{i}]$ into regions of high probability given $Y_{i}=y_{i}$ by the guided distribution introduced in Section 2.2. For this, define
@@
The measure $\mathcal{L}^{g_{i}}(X_{[t_{i-1},t_{i}]})$ is the ‘one-step-ahead’ guided distribution on $t_{i-1},t_{i}]$ corresponding to the next observation $y_{i}$. Hence, the functions $g_{i}$ and $G_{i}$ only need to be computed as functions of $y_{i}$ and the parametrisation of Theorem 2.6 remains a computationally valid choice. Specifically, on any $t_{i-1},t_{i}]$, $g_{i}(t,x)$ is the Gaussian likelihood $g_{i}(t,x)=f(y_{i};L_{t}x+\alpha_{t},R_{t})$ with $L_{t}=LS_{t_{i}-t}$, $\alpha_{t}=L\left(\int_{t}^{t_{i}}S_{t_{i}-s}a_{s}\,\mathop{}\!\mathrm{d}s\right)$ and $R_{t}=\Sigma+LQ_{t_{i}-t}L^{*}$, whereas $G_{i}$ is the score